LBBG Commissioning with AGW1: 11 December 2011 UT

Observer: JHill, JHeidt, KBoutsia (LBTO) DMiller (Tucson)
Telescope Operator: SAllanson (LBTO)
Telescope Support: JUrban (LBTO)
Instrument Support: TShih (LBTO)

Summary

A failed elevation power supply cost us some time at the beginning. As did a failed latch pin on the Right Shutter Door.

We were able to measure the Rotator Center on the AGW1 guider and adjust the probe coordinates to match the hotspot for WFS. Note: The hotspot may still need a slight adjustment to be well centered in the lenslets.

We adjusted the M3 Tip-Tilt to reduce the wobble of the pupil position on the WFS. We got it down to the point where we were having trouble with the measurement (Doug's eyeball) - less than 2 pixels.

We took a set of transform data on Stone D in poor conditions (1.5 arcsec seeing). The preliminary result is that things seem better than Nov 4, but the data has alot (2") of scatter.

We also measured the Rotator Center on IRTC.

We were closed by clouds/humidity at 07:52 UT.

Preparations

The AGW1 WFS camera is working now. We didn't do anything to it other than dismounting the AGW from the telescope.

The SX Adaptive Secondary has been remounted on the telescope, and the SX collimation and pointing models have been returned to those determined on 20111018 with the adaptive secondary. The X-Y-Z global offsets on M2 have been adjusted to remove the coma and focus induced by the remounting of the adaptive secondary.

Jesper has sent us a new transformation which now seems to match the data from 20111103 and 20111104. John has installed oacontrol.conf.20111207b and Steve is restarting oacserver (at 8:25 UT on Dec 10).
From: Jesper Storm 
Date: Wed, 7 Dec 2011 17:09:01 +0100

  I have left modified oacontrol.conf files based on the most recent
oacontrol.conf file, which also contains the new encoder zero point for AGW-3.

The file /lbt/oacontrol/etc/oacontrol.conf.20111207a has the nominal transform
for AGW-1 and oacontrol.conf.20111207b has the most recent fit to the Nov.4
data using the 270 degree relative rotation of the IRTC data.

  You may install your preferred file at your leisure.

Details

23:46 Elevation Drive suffers a power supply failure (or something like that).

Vent doors 1/4 open to cool down the chamber after sunset.

01:30 Back in operation after Jeff and Tim swap in the spare power supply.

Open

01:40 Open to mostly clear, but not photometric sky. T=-3 degC. D=-7.5 degC

Latch Pin 1005 overcurrents while retracting, but we don't need the Right shutter door open. Jeff says the drive is broken on that latch.

01:45 ACQUIRE preset to

ptlist not working ??, cant' change CA from the GUI either. PCS is likely hung because of all the restarting of MCSPU and MCS.

01:50 Kill and restart PCS.

01:51 ACQUIRE preset - but shell won't flatten. Runa is here.

01:57 ACQUIRE preset succeeds, pointing similar to last night. IE=-19 CA=-27.

02:00 ACTIVE preset to BS star on-axis - needs several thousand negative Z6.

Measure Rotator Center on AGW

Stop Guiding

02:09 readGuideCam left -e 72000 guider 0002 is only a partial circle

02:12 readGuideCam left -e 90000 - guider 0003 217,217

02:14 ACTIVE preset off-axis

hotspot coords in AO_IRTC_L.cfg are set to 220,229 (some further adjustment might be needed)

(trying to follow what I did with Jesper's help on 20111101)

02:27 ACTIVE preset on-axis

getxy -u 1 - Guide probe position x: -0.030000 y: 612.472000

setxy -u 1 -x 0 -y 612.8 (increasing y by 330 microns)

stop Guiding

02:33 readGuideCam left -e 92000 - guider 0004 - the very small circle is at 220,230 so this was the correct Y adjustment.

getxy -u 1 - Guide probe position x: 0.088000 y: 612.800000 (millimeters)

getxy -o 16 -u 1 - Guide probe position x: 5.000000 y: 1011997.000000 (encoder counts)

vi oacontrol.conf
         # initial ooffset=417485,&1Q37
                        # adjusted oofset by +2000 microns to 417185 20111101 JMH
                        # adjusted oofset by -330 microns to 416855 20111211 JMH 
                        ooffset=416855,&1Q37

02:49 stopGCSL, stopAGW, restart oacontrol, startAGW, start GCSL - GCS hangs initializing WFS camera
Steve VNCs to azcamserver and restarts the server windows.
Restart GCSL - OK

02:56 ACTIVE preset on-axis

getxy -u 1 - Guide probe position x: 0.073000 y: 612.472000

getxy -o 16 -u 1 - Guide probe position x: 4.000000 y: 1012007.000000

stop Guiding

03:01 readGuideCam left -e 92000 - guider 0005 - small blob at 222,229

dCA=5 to make a larger circle

03:04 readGuideCam left -e 92000 - guider 0006 - 221,230 - OK - AGW probe adjustment complete

IE=-26 CA=-23

Measure Rotator Center on IRTC

03:22 ACTIVE off-axis

Delayed this task until later.......

M3 Alignment to minimize pupil wobble on WFS

Doug on-line via Skype

03:32 ACTIVE on-axis PA=0

setWFSExpTime left 15000

Doug sees pupil center at 377,202 on wfsc 059

03:38 ACTIVE on-axis PA=90

setWFSExpTime left 15000

Doug sees pupil center at 373.5, 198 on wfsc 075

03:42 ACTIVE on-axis PA=180

no stars in WFS on first image (same as the previous preset, but John didn't change exp time) see Issue 3855.

Doug sees 375.5, 195 on wfsc 086

adjust Tip from +100 to +180 arcsec with M3 global offsets.

wfsc 099 = 378,194.5 (wrong image measured?, so we went on the wrong axis?)

03:53 Tip from 180 to +250 arcsec (pupil not measured)

03:55 Tip returned to +100

03:57 Tip +100 Tilt -500 (abs -815, 380 -25.789)

wfsc 132 = 378,196

04:01 Tip +180 Tilt -400

wfsc 150 = 380,194

04:08 Tip +180 Tilt -700 (abs Tip -735, Tilt +180, -25.789)

wfsc 171 = 382, 194

04:17 Tip +400 Tilt -400 (abs -515, +480)

wfsc 205 = 380,192

04:31 Tip -300 Tilt -400 (abs Tip -1215, Tilt +480) (PSF M3 is truncating the - sign)

wfsc 251 = 376,200.5

04:36 ACTIVE on-axis PA=90 (no star - M3 tilted alot)

04:38 ACQUIRE IE=-12 CA=-48

04:42 ACTIVE on-axis PA=90 (first wfs image before star arrived)

wfsc 267 = 380,201

wfsc 270 = 378.5, ??

04:45 ACTIVE on-axis PA=0

wfsc 279 = 379.5, 195

04:54 ACTIVE on-axis PA=-90 (first wfs image before star arrived)

wfsc xxx = 374,196.5

wfsc 293= 375.5, 196.0

0,-90 similar, 90,180 similar

05:06 move Tip from -300 to -200 arcsec (since it looks like we overshot)

05:08 ACTIVE on-axis PA=0

05:11 ACTIVE on-axis PA=90

05:17 ACTIVE on-axis PA=180

Based on these last 4 positions, Doug thinks the total wobble of the pupil position is 2 pixels or less.

The adopted M3 position is therefore dTip = -200, dTilt = -400 which gives the new absolute tertiary position of Tip=-1115 Tilt=+480 Piston 0 RZ=-25.789. This has been installed in oss.conf. SX_LBBG.20111211.ptmod has special IE/CA IE=-18 CA=-45 for this tertiary position.

Transform Data on Stone D

05:27 ACTIVE on-axis Stone D

05:34 start idl after Doug corrected DD problems
  • transform_collect, /GCS -> select Stone_D_new.list
  • the script aborted
 
===> Error in irc_getirparams
        err = Resmsg[0]  = IRTC2 GetIRParams result status: ERROR
        err = Resmsg[1]  = [Error] IRSocket::GetParameters -> IR System busy. Command has been understood, but NOT executed.
% Program caused arithmetic error: Floating illegal operand
  • continous mode was clicked on IRTC - we set it to grab and resend the same command

05:39 transform_collect, /GCS

05:53 preset cancelled because guide probe could not reach the star
  • the GS chosen (27 at PA=0) was on the edge for this PA
  • we chose another GS (25) to replace this and we restarted transform_collect

05:54 transform_collect, /GCS, START=5
  • before sending this routine we changed the irtc image prefix from irtc_20111211 to irtc

06:02 it gave problem also with star GS=41
  • replaced with GS=43

06:04 we started it again after changing some extreme-position stars
  • transform_collect, /GCS, START=7
  • actually this time it did 43 and it failed at 31... but in the list I changed the 43...
  • it will fail again at 31...

06:27 it failed again with GS=41 and PA=-45
  • this time the loaded list should have mostly reachable stars...
  • transform_collect, /GCS, START=12

06:43 failed again trying to reach GS=36 at PA=45
  • it should be reachable....
  • we replaced 36 with 13 and we repeat routine
  • transform_collect, /GCS, START=16

07:11 seeing has been variable - bad at the beginning (1.3''-1.5'') and better at the end (0.9''), clouds are increasing. Humidity at 87%.

07:12 the script failed on the last star (GS=42/PA=90) but we have 28 stars already and we consider the task done

Transform LOG

GStar wfsc # Single Image Cube Image
0 0 irtc.20111211.054005.fits irtc_20111211.20111211.00001_cube.fits
22 360 irtc.20111211.054111.fits irtc_20111211.20111211.00002_cube.fits
30 361-363 irtc.20111211.054240.fits irtc_20111211.20111211.00003_cube.fits
4 364-367 irtc.20111211.054526.fits irtc_20111211.20111211.00004_cube.fits
36 368-369 irtc.20111211.054851.fits irtc_20111211.20111211.00005_cube.fits
GStar wfsc # Single Image Cube Image
25 374 irtc.20111211.055754.fits irtc.20111211.00007_cube.fits
9 375-377 irtc.20111211.055954.fits irtc.20111211.00008_cube.fits
GStar wfsc # Single Image Cube Image
43 381 irtc.20111211.060714.fits irtc.20111211.00010_cube.fits
GStar wfsc # Single Image Cube Image
42 388 irtc.20111211.061356.fits irtc.20111211.00012_cube.fits
GStar wfsc # Single Image Cube Image
10 391 irtc.20111211.061831.fits irtc.20111211.00014_cube.fits
28 392-393 irtc.20111211.061959.fits irtc.20111211.00015_cube.fits
1 394-397 irtc.20111211.062206.fits irtc.20111211.00016_cube.fits
24 398 irtc.20111211.062550.fits irtc.20111211.00017_cube.fits
GStar wfsc # Single Image Cube Image
22 404 irtc.20111211.063146.fits irtc.20111211.00019_cube.fits
11 405-406 irtc.20111211.063315.fits irtc.20111211.00020_cube.fits
25 407-409 irtc.20111211.063522.fits irtc.20111211.00021_cube.fits
2 410-414 irtc.20111211.063806.fits irtc.20111211.00022_cube.fits
GStar wfsc # Single Image Cube Image
13 418 irtc.20111211.064659.fits irtc.20111211.00024_cube.fits
22 419-420 irtc.20111211.064829.fits irtc.20111211.00025_cube.fits
4 421-423 irtc.20111211.065036.fits irtc.20111211.00026_cube.fits
28 424-425 irtc.20111211.065322.fits irtc.20111211.00027_cube.fits
5 426-427 irtc.20111211.065529.fits irtc.20111211.00028_cube.fits
30 428-430 irtc.20111211.065757.fits irtc.20111211.00029_cube.fits
11 431-432 irtc.20111211.070044.fits irtc.20111211.00030_cube.fits
41 433-435 irtc.20111211.070252.fits irtc.20111211.00031_cube.fits
2 436-437 irtc.20111211.070539.fits irtc.20111211.00032_cube.fits
4 438-439 irtc.20111211.070818.fits irtc.20111211.00033_cube.fits

Measure IRTC rotator center

07:19 we started the procedure for measuring IRTC rotation center
  • from iraf : newpoint BS9116
  • it aborted because of rotator fault
Resmsg[0]  = PresetTelescope result status: Error
Resmsg[1]  = Stop[252] complete
Resmsg[2]  = PCS failed StartTrackingRotator on side left because ERROR rotator is not movable (Not 'READY' for some reason)..
Resmsg[3]  = PCS StartTracking failed on side left with DEFAULT.

07:21 Preset to BS9116, gs=0 active, let collimate

07:25 Preset again to BS9116, PA=0 this time track
  • takeseq iraf command has failed because it could not find some DD variables
  • we take the IRTC cube with the idl routines

07:29 irc_getseqimage, 120
  • actually this cube has no source on it...
  • the exposure times on the cube header do not make sense irtc.20111211.00036_cube.fits
  • we repeat the procedure putting manually the exptime

07:34 irc_getseqimage, 120, exp=1
  • now it seems to work fine.
  • the irtc cube for the IRTC rotation center is irtc.20111211.00037_cube.fits
  • again we see nothing - probably the star is too faint, we do have a lot of clouds at this moment

07:39 slew to WT10_233 star to repeat procedure (through heavier clouds)

07:45 irc_getseqimage, 600, exp=0.5
  • irtc.20111211.00038_cube.fits - the cube image to use
  • we tried to create a coadded version of the cube with the idl routine irtc_stack_cube but this is an old routine
  • Jochen does this with MIDAS - it is in /OldRepository/AGW_Data/20111211/final.fits
  • the quality is not great
  • the rotator center we have calculated is: x=158+-1, y=129+-1, which can be considered a first approximation

Close

07:52 close because of clouds and rising humidity

-- DougMiller - 07 Dec 2011
Topic revision: r12 - 11 Dec 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback